OpenStudy (anonymous):

if the parent function is y= 1/x describe the change in the equation y=1/x-7

OpenStudy (anonymous):

does it move up 7 units, move down 7 units, move 7 units left, or 7 units right?

OpenStudy (campbell_st):

it curve moves right 7 units

OpenStudy (anonymous):

what if the chafe in the equation is y=1/(x) -2 ? @campbell_st

OpenStudy (anonymous):

change**

OpenStudy (campbell_st):

always look at the value of x that make the denominator zero... y=1/(x-2) has an asymptote at x = 2 so a shift right... use https://www.desmos.com/calculator to check things

OpenStudy (campbell_st):

but if its \[y=\frac{1}{x} - 2\] the asymptote is at x = 0 and the curve moves down 2

OpenStudy (campbell_st):

to avoid confusion... graph the parent and then graph your choices...